The following reaction is performed: 3H₂ + N₂ → 2NH₃ . You begin this reaction with 10.00 grams of each reactant. What is the excess reactant?
1
Expert's answer
2020-04-23T13:06:49-0400

3H₂ + N₂ → 2NH₃

nH2 = mH2/MrH2 = 10 g / 2 (g/mol) = 5 mol

Per 1 mol: 5/3 = 1,66 mol of H2

nN2 = mN2/MrN2 = 10 g / 28 (g/mol) = 0,357 mol. Its per 1 mol.

So, the excess reactant is N2.